one number is 8 times a first number, a third number is 100 more than the first number, if the sum of all three numbers is 620 then whats the numbers
Answer by rfer(16322)

first=x=52
second=8x=416
third=x+100=152
----------------
x+8x+(x+100)=620
10x=520
x=52